OpenStudy (anonymous):

Will give a medal!! A cyclist must travel 800 km. how many days will the trip take if the cyclist travels 8 h/day at an average speed of 16 km/h?

OpenStudy (anonymous):

About 6 days Write: v=d/t t=d/v t=(800km)/(16km/h) t=50h Since the cyclist travels 8h/day for 50h, we have: 50/8 =6.25 days. or 6days and 1/4 day or 6days and 6hours